# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v2

@@ Appendix C Proof of Theorem 1
[…]
@@
With perfect recovery of the $z^{(i)}[n]$, the terms $|z^{(i)}[n]-\hat{z}^{(i)}[n]|=0$ for all $i$ and $n$. Thus, the MSE is solely due to the quantization noise power after filtering. From Section III-B, the quantization noise power is $\mathbb{E}\{|\epsilon[n]|^{2}\}=\frac{1}{4}(\frac{2\lambda}{2^{b}})^{2}$. This quantization noise power is spread evenly over the entire bandwidth ($-\pi$,$\pi$) since $\epsilon[n]$ is a white process. Considering the filtering operation via a digital LPF with passband region $\left(-\frac{\pi}{\mathrm{OF}}-\Omega_{\mathrm{SL}},+\frac{\pi}{\mathrm{OF}}+\Omega_{\mathrm{SL}}\right)$ after the unfolding step, the power of the filtered quantization noise is only $\frac{1+\frac{K_{\mathrm{SL}}}{N}\cdot\mathrm{OF}}{\mathrm{OF}}$ of $\mathbb{E}\{|\epsilon^{(i)}[n]|^{2}\}$:
$\mathbb{E}\left\{\big{|}\epsilon^{(i)}_{\mathrm{LPF}}[n]\big{|}^{2}\right\}=$ $\frac{\mathbb{E}\{|\epsilon^{(i)}[n]|^{2}\}}{\mathrm{OF}}\left(1+\frac{K_{\mathrm{SL}}}{N}\cdot\mathrm{OF}\right)$
$=$ $\frac{\lambda^{2}}{\mathrm{OF}\cdot 2^{2b}}\left(1+\frac{K_{\mathrm{SL}}}{N}\cdot\mathrm{OF}\right)$
$=$ $\frac{(\lambda^{\prime})^{2}}{\mathrm{OF}\cdot(2^{b}-2)^{2}}\left(1+\frac{K_{\mathrm{SL}}}{N}\cdot\mathrm{OF}\right).$

### v3

@@ Appendix C Proof of Theorem 1
Suppose we let $M=\underset{i}{\max}\;\|\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}^{c}}\|_{\infty}$. The first-order difference of the pre-estimate $\underline{\tilde{z}^{(i)}}[n]$ is at most $\frac{6\lambda^{\prime}}{2^{b}-2}\left(1+M\right)$ away from the first-order difference of the true residue sample. Since a rounding operation to the nearest integer multiple of $2\lambda^{\prime}$ is applied to the $\underline{\tilde{z}^{(i)}}[n]$ to get $\underline{\hat{z}^{(i)}}[n]$, perfect recovery of $z^{(i)}[n]$ after the application of the rounding operation and equation (9) is guaranteed if
@@
With perfect recovery of the $z^{(i)}[n]$, the terms $|z^{(i)}[n]-\hat{z}^{(i)}[n]|=0$ for all $i$ and $n$. Thus, the MSE is solely due to the quantization noise power after filtering. From Section III-B, the quantization noise power is $\mathbb{E}\{|\epsilon[n]|^{2}\}=\frac{1}{4}(\frac{2\lambda}{2^{b}})^{2}$. This quantization noise power is spread evenly over the entire bandwidth ($\pi$$\pi$) since $\epsilon[n]$ is a white process. Considering the filtering operation via a digital LPF with passband region $\left(-\frac{\pi}{\mathrm{OF}}-\delta_{\mathrm{SL}},+\frac{\pi}{\mathrm{OF}}+\delta_{\mathrm{SL}}\right)$ after the unfolding step, the power of the filtered quantization noise is only $\frac{1}{\mathrm{OF}}+\frac{\delta_{\mathrm{SL}}}{\pi}$ of $\mathbb{E}\{|\epsilon^{(i)}[n]|^{2}\}$:
$\mathbb{E}\left\{\big{|}\epsilon^{(i)}_{\mathrm{LPF}}[n]\big{|}^{2}\right\}=$ $\frac{\mathbb{E}\{|\epsilon^{(i)}[n]|^{2}\}}{\mathrm{OF}}\left(1+\frac{\delta_{\mathrm{SL}}\cdot\mathrm{OF}}{\pi}\right)$
$=$ $\frac{\lambda^{2}}{\mathrm{OF}\cdot 2^{2b}}\left(1+\frac{\delta_{\mathrm{SL}}\cdot\mathrm{OF}}{\pi}\right)$
$=$ $\frac{(\lambda^{\prime})^{2}}{\mathrm{OF}\cdot(2^{b}-2)^{2}}\left(1+\frac{\delta_{\mathrm{SL}}\cdot\mathrm{OF}}{\pi}\right).$